# MessIRve: A Large-Scale Spanish Information Retrieval Dataset

### v1

@@ Appendix A Supporting information
Table 3 shows the seed prefixes used to obtain the queries for each country in the MessIRve dataset. These are loosely based on Khashabi et al. (2021). We include incorrect, non-accented, versions of words, e.g. “como” instead of “cómo”, to account for common spelling mistakes in Spanish queries.
@@
Table 4 shows the most distinctive words for each country in the MessIRve dataset. These words are obtained by calculating the log odds ratio of the frequency of each word in the queries for each country with respect to the queries of the full subset, and selecting the words with the highest scores. The words are indicative of the dialectal variety of Spanish spoken in each country.
@@
Table 5 shows the most representative words and queries for the 8 most frequent topics in each dataset. Representative words and queries were selected based on the c-TF-IDF scores (Grootendorst, 2022).
@@
Table 4: Words in the queries of MessIRve are representative of the dialectal varieties of Spanish and the topics of interest in each country. To find distinctive words for each country, we calculated the log odds ratio of the frequency of each word in each country with respect to the queries of the full subset, using smoothing with a prior of 0.1 (Jurafsky and Martin, 2024). We consider only words with at least 10 occurrences in the country subset and show the top 6 words with the highest log odds ratio for each country.

### v2

@@ Appendix C Queries
Table 3 shows the seed prefixes used to obtain the queries for each country in the MessIRve dataset. These are loosely based on khashabi-etal-2021-gooaq-open. We include incorrect, non-accented, versions of words, e.g. “como” instead of “cómo”, to account for common spelling mistakes in Spanish queries.
The 120 prefixes used are seed prefixes. For each country, we send each seed prefix to the API with all one-letter continuations (1+27 API calls per prefix). We compile results, retain those starting with a seed prefix, and generate new prefixes adding 1 word. Then we consider 2-word continuations, and stop when reaching a target results count per country. The number of API results per call varies: more specific prefixes often yield fewer results. We do not extract a fixed number of completions per prefix, and we retain queries not starting with seed prefixes (but don’t use them for further API calls).
@@
Table 4 shows the most distinctive words for each country in the MessIRve queries. These words are obtained by calculating the log odds ratio of the frequency of each word in the queries for each country with respect to the queries of the full subset, and selecting the words with the highest scores. The words are indicative of the dialectal variety of Spanish spoken in each country.
@@
Table 5 shows the most representative words and queries for the 8 most frequent topics in each dataset. Representative words and queries were selected based on the c-TF-IDF scores (grootendorst2022bertopic).
Representative query-document pairs per country are shown in Table 6. Queries are sampled from those containing the top log-odds ratio words per country, as listed in Table 4.
MessIRve contains shorter queries than other datasets, but with higher prefix diversity. This is shown in Figure 1 (distribution of query lengths in each dataset) and Figure 5 (entropy and number of unique prefixes of length 1 to 6 in each dataset).
@@
Table 4: Words in the queries of MessIRve are representative of the dialectal varieties of Spanish and the topics of interest in each country. To find distinctive words for each country, we calculated the log odds ratio of the frequency of each word in each country with respect to the queries of the full subset, using smoothing with a prior of 0.1 (jurafsky2024slp). We consider only words with at least 10 occurrences in the country subset and show the top 6 words with the highest log odds ratio for each country.
@@
Country Query Relevant document
ar quien era mama antula (who was Mama Antula?) María Antonia de Paz y Figueroa. Mama Antula se convirtió en la novena persona de nacionalidad argentina en ser beatificada. Esto ocurrió entre casi medio centenar de causas para canonizar […]
bo a cuantos kilometros esta samaipata de santa cruz (how many kilometers is Samaipata from Santa Cruz?) Samaipata. La ciudad de Samaipata se encuentra a 119 km por carretera al suroeste de la capital departamental, Santa Cruz de la Sierra. Por Samaipata discurre la ruta troncal Ruta […]
cl en qué región queda huilo huilo (in which region is Huilo Huilo located?) Reserva biológica Huilo Huilo. La Reserva Biológica Huilo Huilo es un área natural protegida privada que se ubica en medio de Los Andes, a 860 km al sur de Santiago […]
co donde murio jorge eliecer gaitan (where did Jorge Eliécer Gaitán die?) Jorge Eliécer Gaitán. Un hombre (aparentemente Juan Roa Sierra u otros más) lo esperaba en la entrada del edificio y le disparó con un revólver causándole heridas mortales. Gaitán fue […]
cr que visitar en puntarenas (what to visit in Puntarenas?) Puntarenas (ciudad). La ciudad de Puntarenas, como capital de la provincia, cuenta con diversos lugares históricos de interés turístico, entre los que se pueden mencionar la antigua Capitanía del Puerto […]
cu que hizo antonio maceo por cuba (what did Antonio Maceo do for Cuba?) Antonio Maceo. José Antonio de la Caridad Maceo y Grajales (San Luis, Santiago de Cuba, 14 de junio de 1845 - San Pedro, La Habana, 7 de diciembre de 1896) […]
do qué hizo enriquillo (what did Enriquillo do?) Enriquillo. Enrique Bejo (lago Jaragua, Cacicazgo de Jaragua, ca. 1498-Sabana Buey, Llano de Baní, 27 de septiembre de 1535), más conocido como Enriquillo, fue un noble taíno que se alzó […]
ec de que region son los montubios (from which region are the Montubios?) Montuvio. Según el Censo ecuatoriano de 2022 los montuvios representan el 7,7% de la población del Ecuador, lo que significa que en el 2022 más de 1.304.994 ecuatorianos se identificaron […]
es de qué se pueden hacer las croquetas (what can croquettes be made from?) Croqueta. A veces, pero no siempre, se enharinan primero y se sacude la harina sobrante. Se bañan en huevo batido y se cubren siempre generosamente con pan rallado. Después se […]
gt dónde queda mazatenango (where is Mazatenango?) Mazatenango. Mazatenango (del náhuatl, significa «"muralla del venado"») es una ciudad y cabecera del departamento de Suchitepéquez, en la República de Guatemala. El municipio se encuentra localizado a 161 km […]
hn quien y en que año se fundo la villa de san pedro sula (who and in which year was the town of San Pedro Sula founded?) San Pedro Sula. San Pedro Sula fue fundada el 27 de junio de 1536, bajo el nombre de San Pedro de Puerto Caballos, por el conquistador español Pedro de Alvarado. […]
mx maluma carin leon - según quién versuri romana (Maluma Carin León - according to whom, Roman lyrics?) Según quién. «Según quién» es una canción del cantante colombiano Maluma y el cantante mexicano Carín León. Fue lanzada el 17 de agosto de 2023, a través de Sony Music […]
ni cuándo hizo erupción el volcán masaya (when did Masaya Volcano erupt?) Volcán Masaya. Otras erupciones han ocurrido en los últimos 50 años. El 22 de noviembre de 1999 empezó un nuevo evento eruptivo, apareciendo un punto rojo en las imágenes de […]
pa que año nacio delfia cortez (in what year was Delfia Cortez born?) Delfia Cortez. Delfia Nereida Cortez Marciaga (Los Santos, 21 de enero de 1963-Ciudad de Panamá, 10 de enero de 2024), fue una periodista y política panameña.
pe quién es dina boluarte (who is Dina Boluarte?) Dina Boluarte. Fue electa primera vicepresidenta en las elecciones presidenciales de Perú de 2021, y ocupó el cargo desde el 28 de julio de 2021 hasta el 7 de diciembre […]
pr que hace el decadron inyectable (what does injectable Decadron do?) Dexametasona. La dexametasona es un potente glucocorticoide sintético con acciones que se asemejan a las de las hormonas esteroides. Actúa como antiinflamatorio e inmunosupresor. Su potencia de 20-30 veces la […]
py cómo surgió la nación paraguaya (how did the Paraguayan nation arise?) Historia de Paraguay. Lo que actualmente es el territorio de Paraguay fue descubierto por Alejo García y Juan de Ayolas, a las órdenes de España en 1524, dándose inicio a […]
Table 6: Example query-document pairs per country. Queries are sampled from those containing the top log-odds ratio words per country, as listed in Table 4. For readability, queries were translated into English by the author, and only the first 30 words of each document are displayed.